OpenStudy (anonymous):

x/5 - y/10 = 4 and 3x/4 + 2y/3 = 8 Use substitution to find the solutions (if any).

OpenStudy (vijay):

x = 416/25 and y = -168/25

OpenStudy (anonymous):

Okay. So what you are being asked to do is this. Solve one equation and then plug in the value of the variable into the other. So starting with the first equation: (x/5 - y/10 = 4) Multiply both sides by 5. Then you have; x - y/2 = 4 Multiply both sides by 2: 2x - y = 4 -y = -2x + 4 y = 2x - 4 Now we can substitute for y in the second equation (3x/4 + 2y/3 = 8) 3x/4 + 2(2x-4)/3 = 8 3x/4 + (4x - 8)/3 = 8 Let's eliminate the denominators: Multiply both sides by 4. 3x + 4(4x-8)/3 = 8 Multiply both sides by 3. 9x + 4(4x-8) = 8 Distribute: 9x + 16x - 32 = 8 Add like terms and add 32 to both sides. 25x = 24 x = 24/25 y = 2x - 4 There you go. If this helps, please leave a medal. Check out my website, where there are free math videos at http://www.tutorsean.net